# Benchmarking LLMs for Translating Classical Chinese Poetry: Evaluating Adequacy, Fluency, and Elegance

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable performance in general translation tasks. However, the increasing demand for high-quality translations that are not only adequate but also fluent and elegant. To assess the extent to which current LLMs can meet these demands, we introduce a suitable benchmark for translating classical Chinese poetry into English. This task requires not only adequacy in translating culturally and historically significant content but also a strict adherence to linguistic fluency and poetic elegance. Our study reveals that existing LLMs fall short of this task. To address these issues, we propose RAT, a Retrieval-Augmented machine Translation method that enhances the translation process by incorporating knowledge related to classical poetry. Additionally, we propose an automatic evaluation metric based on GPT-4, which better assesses translation quality in terms of adequacy, fluency, and elegance, overcoming the limitations of traditional metrics. Our dataset and code will be made available.
Benchmarking LLMs for Translating Classical Chinese Poetry:
Evaluating Adequacy, Fluency, and Elegance
Andong Chenthanks: Work was done when Andong Chen was at Pengcheng Laboratory. , Lianzhang Lou , Kehai Chenthanks:   Corresponding author. , Xuefeng Bai , Yang Xiang , Muyun Yang , Tiejun Zhao , Min Zhang School of Computer Science and Technology, Harbin Institute of Technology, China Pengcheng Laboratory, Shenzhen, China ands691119@gmail.com, {loulzh, xiangy}@pcl.ac.cn {chenkehai, baixuefeng, yangmuyun, tjzhao, zhangmin2021}@hit.edu.cn,
@@ 1 Introduction
The emergence of large language models (LLMs), especially ChatGPT, has demonstrated impressive performance in general translation tasks Tyen et al. (2023); Liang et al. (2023); Guerreiro et al. (2023); Ranaldi et al. (2023). The extensive knowledge and strong reasoning abilities of these LLMs open new possibilities in the field of translation Ouyang et al. (2022); Moslem et al. (2023b); Chen et al. (2023); He et al. (2023); Liang et al. (2023); Ghazvininejad et al. (2023); Lu et al. (2023); Chen et al. (2024a). As the demand for translation quality continues to rise, translated results must not only be adequate but also fluent and elegant Wang et al. (2024); Huang et al. (2024); Gao et al. (2024). This raises a question: can existing LLMs meet such translation requirements, and if so, to what extent can they achieve this performance?
To address this question, we introduced a suitable benchmark: translating classical Chinese poetry into English. Firstly, these poems carry culture and history, so the translation needs to adequately convey the original meaning. Secondly, classical Chinese poetry has strict rules on rhyme, tone, and structure, making fluent translation a significant challenge. Lastly, classical Chinese poetry has aesthetic value, with the concise expressions of the Chinese language showing unique elegance, which needs to be preserved in translation.
Based on the proposed classical poetry translation benchmark, current methods often lack historical and cultural knowledge, fluency, and elegance. To address these issues, we introduce RAT, a retrieval-augmented machine translation method. This method enhances the translation process by retrieving knowledge related to classical poetry, ensuring the translated output remains adequate to the original text while also achieving fluency and elegance. Furthermore, previous automatic evaluation metrics for machine translation Papineni et al. (2002); Rei et al. (2022); Sellam et al. (2020a) have primarily relied on n-gram and semantic analysis to assess entire sentences, making them unsuitable for the task of classical poetry translation. To address this issue, we propose an automatic evaluation metric based on GPT-4 (OpenAI 2024), which better assesses translation quality from the perspectives of adequacy, fluency, and elegance.
To our knowledge, this is the first evaluation and research on the translation performance of LLMs based on the task of translating classical Chinese poetry. Through this effort, we aim not only to test the capabilities of LLMs in translating classical Chinese poetry but also to inspire community discussion on the potential and future development of large models in translating texts that are adequate, fluent, and elegant.
@@
  • We have introduced the first classical poetry translation dataset, which allows for better evaluation of LLMs in terms of adequacy, fluency, and elegance.
  • Based on the limitations of current LLM-based translation methods on the classical poetry translation dataset, we have proposed a retrieval-augmented translation method to further enhance the performance of large models in this task.
  • We have designed a new evaluation metric based on GPT-4 to assess classical poetry translation. This metric aligns more closely with human annotations and is better suited for evaluating the translation of classical poetry.
@@ 2 Related Work
Poetry Machine Translation The initial attempts at poetry machine translation utilized phrase-based machine translation systems to translate French poetry into metrical English poetry Genzel et al. (2010). This study demonstrated how statistical machine translation systems can produce translations while adhering to the rhythmic and rhyming rules of poetry. Chakrabarty et al. (2021) conducted an empirical study that highlighted a critical but often overlooked issue: even though they can maintain meaning and fluency, advanced machine translation systems trained on large amounts of non-poetry data struggle to preserve the original style of poetry. To address the issue of style preservation, some studies have managed to generate diverse styles by inputting sentences with specific styles into the encoder and embedding the target style into the decoder Zhang et al. (2018); Liu and Wang (2012). Considering that poetry often contains richly stylized semantic information and deep historical and cultural connotations, such as in classical Chinese poetry, Rajesh Kumar Chakrawarti and Bansal (2022) proposed a Hybrid Machine Translation (HBMT) model to improve the semantic and syntactic accuracy of poetry translations. Recently, Wang et al. (2024) have successfully implemented translations of modern poems from English to Chinese using the knowledge and multilingual capabilities of ChatGPT.
Chinese Classical Poetry Dataset Chinese classical poetry has several open datasets. Chen et al. (2019) published the first fine-grained dataset, annotating a manually fine-grained emotional poetry corpus containing 5,000 Chinese quatrains. Yutong et al. (2020) released a dataset containing 3,940 quatrains with automatically annotated themes and 1,917 quatrains with automatically annotated emotions, using a template-based method. Additionally, Liu et al. (2020) collected a parallel bilingual dataset of ancient and modern Chinese and aligned the lines using a string matching algorithm. Based on these bilingual pairs, Li et al. (2021) constructed a matching dataset to evaluate models’ semantic understanding capabilities. To our knowledge, this is the first dataset for English translation of Chinese classical poetry, aimed at evaluating the translation performance of current large models in terms of “adequate, fluent, and elegant".
@@ 3 Classical Chinese Poetry Dataset Construction
In this section, we discuss the design and construction of the MT benchmark, including the rules and steps for building this benchmark.
@@ 3.1 Discourse-Level Poetry Translation
We collected a batch of classical Chinese poetry data and corresponding English translations from online resources. The classical Chinese poetry database contains 1200 poems from Tang Poems, Song Poems, and Yuan Opera. Then, We manually screened 608 classical Chinese poems and their corresponding translations to serve as gold standards for translation source and target alignment. As shown in Figure 1.
The statistics of this benchmark are shown in Table 1. We present the number of classical Chinese poetry, the number of unique tokens, the average number of tokens per sentence, and the total number of tokens in different poetry types. The source sentences in this benchmark have a moderate length, and the selected target translation sentences are well-aligned with the source in terms of length, indirectly reflecting the high quality of the reference sentences.
@@
Table 1: Statistics on the benchmark. Numbers a/b denote the corresponding number in source sentences/target sentences.
@@ 3.2 Adequacy in Sentence-Level Translation
Figure 2: An example block in the elegance in discourse-Level poetry translation.
To construct this test set, we selected sentences containing historical knowledge and commonsense from the collected data of classical Chinese poetry. To maintain the diversity of the test set, we avoided choosing semantically similar words and those that, despite having multiple meanings in Chinese, translate to the same word in English. We preferred to select words that translate to completely different English words in different contexts, requiring disambiguation based on understanding various contexts, historical backgrounds, writing styles, and commonsense. This test set consists of 500 sentences, with each translated sentence represented as a triplet ($s$, $t_{c}$, $t_{e}$). As shown in Figure 2, $s$ contains the source sentence with ambiguous words, $t_{c}$ and $t_{e}$ are the contrast translations, with the former being correct and the latter incorrect. The selected ambiguous words are used only once in each translated sentence.
@@ 3.3 Classical Chinese Poetry Knowledge Base
Classical Chinese Poetry holds rich historical and cultural nuances, but with limited resources for Classical Chinese, modern Chinese knowledge can greatly improve translation quality. We collected a Classical Chinese Poetry Knowledge Base using open-source projects and internet resources. This Knowledge Base consists of 30,000 entries, including 30,000 Classical Chinese Poems along with their corresponding historical background, dynasty names, modern Chinese translations, author introductions, modern Chinese analysis, and poetry type information.
3.4 Human Annotation
Although collected poetry translation data were created by humans, the randomness still affects data quality. To mitigate this, we invite human annotators to assess the quality of translation parallel sentence pairs and eliminate any problematic instances.
@@ 4 Proposed Method: RAT
Figure 3: The proposed RAT framework.
Translation of classical Chinese poetry requires understanding of language style, cultural background, and author information. The RAT (Retrieval Augmented Translate) method was introduced to enhance translation outcomes by leveraging rich contextual information from the Classical Chinese Poetry Knowledge Base. We will now introduce the two workflows implemented in the RAT method in detail (in Figure 1). The first workflow (in Figure 1(a)) uses poetry as input and retrieves relevant knowledge about the poetry from the Classical Chinese Poetry Knowledge Base through text matching methods, obtaining different views of the relevant knowledge. The second workflow (in Figure 1(b)) translates based on the relevant knowledge from different views and integrates these to produce the final translation result.
@@ 4.1 The First Workflow
In the first workflow of RAT, there are two modules: Retriever and Selector. The Retriever is used to retrieve relevant knowledge from diverse perspectives based on the poetry, while the Selector filters the retrieved knowledge to enhance its relevance to the poetry.
1. Retriever. We propose a retrieval augmentation method to obtain knowledge relevant to translating classical Chinese poetry. Based on the Classical Chinese Poetry Knowledge Base we constructed, we use text-matching methods to retrieve uniquely relevant knowledge from multiple perspectives. These perspectives include historical background, dynasty name, modern Chinese translation, author introduction, modern Chinese analysis, and poetry type. The retrieved examples are shown in Appendix A.
2. Selector. The goal of the selector is to filter out irrelevant content from the results of the retriever that are unrelated to the translated sentences and passages. As an agent set by the LLM, the selector first understands the historical background, author introduction, and modern Chinese analysis based on the source sentences, and then outputs content relevant to the translated source sentences, with a limit on the output length (max_len). Specific prompts are displayed in Appendix A.1.
@@ 4.2 The Second Workflow
In the second workflow of RAT, there are three modules: Translator, Voter, and Extractor. The Translator module translates different English verses based on classical Chinese poetry and the retrieved knowledge. The Voter module then votes on each translated sentence to determine the optimal translation. Finally, the Extractor module extracts the voted translation results to form the final translated version.
@@
1.Translator. The goal of the Translator is to translate different types of retrieved content. The LLM uses its translation capabilities to process both source content and retrieved content separately. This time, six types of related content were retrieved for classical Chinese poetry, resulting in six different translation outputs. Specific prompts are displayed in Appendix B.1.
@@
2.Voter. The purpose of the Voter is to integrate translations based on different retrieval results to improve translation quality. As an agent of the LLM, Voter votes on each sentence of all translation outputs based on the source input, selects the highest-quality translations, and concatenates these selected sentences to form the final translation result. Specific prompts are displayed in Appendix B.2.

### v4

# Large Language Models for Classical Chinese Poetry Translation: Benchmarking, Evaluating, and Improving

@@ Abstract
Different from the traditional translation tasks, classical Chinese poetry translation requires both adequacy and fluency in translating culturally and historically significant content and linguistic poetic elegance. Large language models (LLMs) with impressive multilingual capabilities may bring a ray of hope to achieve this extreme translation demand. This paper first introduces a suitable benchmark (PoetMT) where each Chinese poetry has a recognized elegant translation. Meanwhile, we propose a new metric based on GPT-4 to evaluate the extent to which current LLMs can meet these demands. Our empirical evaluation reveals that the existing LLMs fall short in the challenging task. Hence, we propose a Retrieval-Augmented machine Translation (RAT) method which incorporates knowledge related to classical poetry for advancing the translation of Chinese Poetry in LLMs. Experimental results show that RAT consistently outperforms all comparison methods regarding wildly used BLEU, COMET, BLEURT, our proposed metric, and human evaluation.
Large Language Models for Classical Chinese Poetry Translation:
Benchmarking, Evaluating, and Improving
Andong Chenthanks: Work was done when Andong Chen was at Pengcheng Laboratory.1 , Lianzhang Lou2 , Kehai Chenthanks:   Corresponding author.1,2 , Xuefeng Bai1 , Yang Xiang2 , Muyun Yang1 , Tiejun Zhao1 , Min Zhang1 1 School of Computer Science and Technology, Harbin Institute of Technology, China 2 Pengcheng Laboratory, Shenzhen, China ands691119@gmail.com, {loulzh, xiangy}@pcl.ac.cn {chenkehai, baixuefeng, yangmuyun, tjzhao, zhangmin2021}@hit.edu.cn,
@@ 1 Introduction
The emergence of large language models (LLMs), especially ChatGPT, has demonstrated impressive performance in translation tasks Tyen et al. (2023); Liang et al. (2023); Guerreiro et al. (2023); Ranaldi et al. (2023); Zhao et al. (2024); Zhang et al. (2024); Chen et al. (2024a). As the requirements for translation quality continues to rise, translated results not only be adequate but also fluent and elegant Wang et al. (2024); Huang et al. (2024); Gao et al. (2024); Wu et al. (2024). This raises a question: can existing LLMs meet such translation requirements, and if so, to what extent can they achieve this performance?
To answer this question, we introduced a suitable benchmark (PoetMT): translating classical Chinese poetry into English. Firstly, these poems carry culture and history, so the translated results need to adequately convey these meanings. Secondly, classical Chinese poetry has strict rules on rhyme, tone, and structure, making fluent translation a significant challenge. Lastly, classical Chinese poetry has aesthetic value, with the concise expressions of the classical Chinese language showing linguistic poetic elegance, which needs to be preserved in translated results.
Based on the proposed PoetMT benchmark, previous automatic evaluation metrics for machine translation only analyze entire sentences without evaluating classical poetry translation quality explicitly Papineni et al. (2002); Rei et al. (2022); sellam-etal-2020-bleurt. To overcome the limitations of traditional evaluation metrics, we propose an automatic evaluation metric based on GPT-4 (OpenAI 2024), which better evaluates translation quality from the perspectives of adequacy, fluency, and elegance. Additionally, evaluating current LLM-based MT methods reveals that these translated results often lack historical and cultural knowledge, strict rhyme and structure rules, and concise expressions. To address these issues, we introduce RAT, a retrieval-augmented machine translation method. This method enhances the translation process by retrieving knowledge related to classical poetry, ensuring the translated output remains adequate to the original text while also achieving fluency and elegance.
To our knowledge, this is the first study evaluating the translation performance of LLMs based on the task of translating classical Chinese poetry. Through this effort, we aim not only to test the capabilities of LLMs in translating classical Chinese poetry but also to inspire community discussion on the potential and future development of LLMs in translated texts that are adequate, fluent, and elegant.
@@
  • We have introduced the first classical poetry translation benchmark (PoetMT), which allows for a better evaluation of LLMs in terms of adequacy, fluency, and elegance.
  • We have designed a new evaluation metric based on GPT-4 to evaluate classical poetry translation. This metric aligns more closely with human annotations and is better suited for the PoetMT benchmark.
  • Based on the limitations of current LLM-based translation methods on the PoetMT benchmark, we have proposed a retrieval-augmented translation method to enhance the performance of LLMs in this task.
@@ 2 Related Work
Poetry Machine Translation The initial attempts at poetry machine translation utilized phrase-based machine translation systems to translate French poetry into metrical English poetry Genzel et al. (2010). This study demonstrated how statistical machine translation systems can produce translations while adhering to the rhythmic and rhyming rules of poetry. chakrabarty-etal-2021-dont conducted an empirical study that highlighted a critical but often overlooked issue: even though they can maintain meaning and fluency, advanced machine translation systems trained on large amounts of non-poetry data struggle to preserve the original style of poetry. To address the issue of style preservation, some studies have managed to generate diverse styles by inputting sentences with specific styles into the encoder and embedding the target style into the decoder Zhang et al. (2018); Liu and Wang (2012). Considering that poetry often contains richly stylized semantic information and deep historical and cultural connotations, such as in classical Chinese poetry, Rajesh Kumar Chakrawarti and Bansal (2022) proposed a Hybrid Machine Translation model to improve the semantic and syntactic accuracy of poetry translations. Recently, Wang et al. (2024) have successfully implemented translations of modern poems from English to Chinese using the knowledge and multilingual capabilities of ChatGPT.
Chinese Classical Poetry Dataset Chinese classical poetry has several open datasets. Chen et al. (2019) published the first fine-grained dataset, annotating a manually fine-grained emotional poetry corpus containing 5,000 Chinese quatrains. Yutong et al. (2020) released a dataset containing 3,940 quatrains with automatically annotated themes and 1,917 quatrains with automatically annotated emotions, using a template-based method. Additionally, Liu et al. (2020) collected a parallel bilingual dataset of ancient and modern Chinese and aligned the lines using a string-matching algorithm. Based on these bilingual pairs, Li et al. (2021) constructed a matching dataset to evaluate models’ semantic understanding capabilities. To our knowledge, this is the first benchmark for English translation of Chinese classical poetry, aimed at evaluating the translation performance of current large models in terms of “adequate, fluent and elegant".
@@ 3 Classical Chinese Poetry Dataset Construction
In this section, we discuss the design and construction of the PoetMT benchmark, including the rules and steps for building this benchmark.
@@ 3.1 Discourse-Level Poetry Translation
Table 1: Statistics on the benchmark. Numbers a/b denote the corresponding number in source/target sentences.
We collected a batch of classical Chinese poetry data and corresponding human English translations from online resources [^1] We selected professional translation versions available online, specifically those authored by Xu Yuanchong. Therefore, the translation results are from very experienced poem translators. The classical Chinese poetry database contains 1200 poems from Tang Poems, Song Poems, and Yuan Opera. Then, We manually screened 608 classical Chinese poems and their corresponding translations to serve as gold standards for translation source and target alignment [^2] The manually filtered content primarily excludes poems with multiple translation versions, misaligned or incomplete Chinese and English poems, and poems with factual errors.. As shown in Figure 1.
The statistics of the PoetMT benchmark are shown in Table 1. We present the number of classical Chinese poetry, the number of unique tokens, the average number of tokens per sentence, and the total number of tokens in different poetry types. The source sentences in this benchmark have a moderate length, and the selected target translation sentences are well-aligned with the source in terms of length, indirectly reflecting the high quality of the reference sentences.
@@ 3.2 Adequacy in Sentence-Level Translation
Figure 2: An example block in the adequacy in sentence-level poetry translation.
Due to the inclusion of historical background and common knowledge in classical Chinese poetry, achieving adequacy in translation poses a significant challenge. Therefore, to conduct a more detailed evaluation of adequacy, we have constructed a sentence-level test set.
Following related worksHe et al. (2020); Yao et al. (2024), we selected sentences containing historical knowledge and commonsense from the collected data of classical Chinese poetry. We avoid selecting semantically similar words to ensure diversity in the test set. Additionally, we preferred to select words that have different English translations depending on the context. This test set consists of 758 sentences, with each translated sentence represented as a triplet ($s$, $t_{c}$, $t_{e}$). As shown in Figure 2, $s$ contains the source sentence with ambiguous words, $t_{c}$ and $t_{e}$ are the contrast translations, with the former being correct and the latter incorrect.
@@ 3.3 Classical Chinese Poetry Knowledge Base
Classical Chinese poetry holds rich historical and cultural nuances, but due to the limited resources for Classical Chinese, modern Chinese knowledge can greatly mitigate this issue. The PoetMT benchmark includes a Classical Chinese Poetry Knowledge Base collected from open-source projects and internet resources. This Knowledge Base consists of 30,000 entries, including 30,000 Classical Chinese poems along with knowledge such as their corresponding historical background, dynasty name, modern Chinese translation, author introduction, modern Chinese analysis, and poetry type. The case is displayed in Appendix C.4.
4 LLM-based Evaluation Method
4.1 Evaluation Criteria
Figure 3: Examples of different evaluation metrics. Figure (a) represents the rhyme of the final words; Figure (b) shows that the two translated sentences have the same word count and couplet structure; Figure (c) indicates the accurate translation capturing the implied meaning of time passing.
The translation of classical poetry requires not only artistic expression but also an understanding of the cultural background, yet the premise of correctness does not imply a singular or unique expression. Following this line of thought, we evaluate the quality of classical poetry translations through three aspects: “Adequate," “Fluent," and “Elegant."
4.1.1 Adequate Criteria
4.1.2 Fluent Criteria
Beauty of Sound (BS)↑: The beauty of sound in Chinese classical poetry is primarily reflected in its rhyme. This standard examines whether the translation achieves harmonious sound, adherence to strict metrical rules, and a rhythm that is both smooth and dynamic. As shown in Figure 3(a).
Beauty of Form (BF)↑: Chinese classical poetry emphasizes symmetrical structures, with common forms including the "Five-character eight-line regulated verse (wulü)", "Seven-character eight-line regulated verse (qilü)", and "Extended forms (pailü)" among others. Each form showcases the structural characteristics of Chinese poetry. This standard evaluates whether the translation maintains consistency with the source poem’s structure, including the alignment of line numbers and balanced phrasing. As shown in Figure 3(b).
4.2 Elegant Criteria
Beauty of Meaning (BM)↑: Chinese classical poetry uses concise and precise language to create vivid imagery and a rich atmosphere for readers. The criteria evaluate the depth and richness of the translation, focusing on the effectiveness of conveying themes, emotions, and messages. As shown in Figure 3(c).
4.3 LLM-based Classical Poetry Metric
In this section, we aim to establish a method for evaluating the translation of classical Chinese poetry based on LLMs. Drawing from current QE (Quality Estimation) research Li et al. (2023); Kocmi and Federmann (2023), we have developed an approach to utilize LLM in determining the quality of classical Chinese poetry translations. We designed a 1-5 scoring prompt to help LLM focus on the translation quality in terms of Beauty of Sound (LLM-BS). In the evaluation criteria, a score of 1 reflects a poor translation of classical Chinese poetry, a score of 3 indicates a basic but imperfect translation result and a score of 5 represents an excellent and accurate translation in this aspect. Subsequently, test data is provided to the LLM, and the LLM is asked to generate an evaluation with scores only. Additionally, we calculate the average scores (LLM-Avg) of BF, BF, and BM to evaluate the overall translation performance. The prompt details for Beauty of Sound(LLM-BS), Beauty of Form (LLM-BF) and Beauty of Meaning (LLM-BM) are provided in Appendix B.6, B.7 and B.8.
@@ 5 Proposed Method: RAT
Figure 4: The proposed RAT framework. The "Historical Background," "Author Introduction," and "Modern Chinese Analysis" parts are at the discourse level, so the Selector needs to make selections based on the content.
The RAT (Retrieval Augmented Translate) method was introduced to enhance translation performance by leveraging rich contextual information from the Classical Chinese Poetry Knowledge Base. We introduce the two workflows implemented in the RAT method in detail (in Figure 4). The first workflow uses poetry as input and retrieves relevant knowledge about the poetry from the Classical Chinese Poetry Knowledge Base through text-matching methods, obtaining different views of the relevant knowledge. The second workflow translates based on the relevant knowledge from different views and integrates these to produce the final translation result.
@@ 5.1 The First Workflow
In the first workflow of RAT, there are two modules: Retriever and Selector.
Retriever. We propose a retrieval augmentation method to obtain knowledge relevant to translating classical Chinese poetry. Based on the Classical Chinese Poetry Knowledge Base, we use text-matching methods to retrieve uniquely relevant knowledge from multiple perspectives [^3] The Classical Chinese Poetry Knowledge Base contains all 608 classical poems presented in the paper, ensuring a one-to-one correspondence between the poems and the knowledge.. These perspectives include historical background, dynasty name, modern Chinese translation, author introduction, modern Chinese analysis, and type.
Selector. The goal of the selector is to filter out irrelevant content to translated sentences from the results of the retriever. As an agent of the LLM, the selector first understands the historical background, author introduction, and modern Chinese analysis based on the source sentences, and then outputs content relevant to the translated source sentences. Specific prompts are displayed in Appendix B.1.
@@ 5.2 The Second Workflow
In the second workflow of RAT, there are three modules: Translator, Voter, and Extractor.
@@
Translator. The goal of the Translator is to translate classical Chinese poetry based on different types of retrieved knowledge. Six types of related knowledge were retrieved for classical Chinese poetry, resulting in six different translation outputs. Specific prompts are displayed in Appendix B.2.
@@
Voter. The purpose of the Voter is to integrate translations based on different retrieval results to improve translation quality. As an agent of the LLM, the Voter votes on each sentence of all translation outputs based on the source input selects the highest-quality translations and concatenates these selected sentences to form the final translation result. Specific prompts are displayed in Appendix B.3.